perform the following conversions using unit factoring.
1. 500mL= L
2. 25cg= g
3. 400mg= kg

Answer by stanbon(75887)

perform the following conversions using unit factoring.
1. 500mL(L/1000mL) = (1/2)L
2. 25cg(g/100cg) = (1/4)g
3. 400mg(g/1000mg)(kg/1000g)= [400/10^6] = (4/10000) = (1/2500)kg
=================
Cheers,
Stan H.